OpenStudy (anonymous):

first derivative of the following: i(x):= 4^x tan(x)

OpenStudy (anonymous):

\[(f*g)'=f'g+fg'\]

OpenStudy (anonymous):

\[f=4^x,g=\tan x\]

OpenStudy (anonymous):

\[f'=4^x \ln 4,g'=\sec ^2x\]